If f(x)=2x2+3a+5, and f(1)=-2, then a = ?
Answer by checkley77(12844)

If f(x)=2x^2+3a+5, and f(1)=-2, then a = ?
f(1)=-2
-2=2*-1^2+3a+5
-2=2*1+3a+5
3a=-2-2-5
3a=-9
a=-9/3
a=-3 ans.